Inlet pipe A,B and C can filled the tank in 12, 15, and 20 hr.if open in order 6, 7, and 9 am. Then the tank will be filled in what time.

Grade:11

3 Answers

Priyanka
6 years ago
In i hour, parts filled by A,B,C tanks are 1/12,1/15,1/20.
 
A,B.C together can fill \frac{1}{12}+\frac{1}{15}+\frac{1}{20}=\frac{1}{5} part of tank in 1 hour.
Thus, 1 tank can be completely filled in 5 hours.
Therefore, The tank will be filled by 11 am. (6 am to 11 am= 5 hours)
Arun
6 years ago
In 1 hour, if three are open then total part filled is1/12 + 1/15 + 1/20 = 1/5 and this will happen only after 9 amHowever Pipe A will be open for additionally 3 hours,Hence part filled = 3/12 = 1/4Pipe B will be open for additionally 2 hours, hence part filled = 2/15Now till 9 am toral part filled = 1/4 + 2/15 = 23/60 Part remained = 1-23/60. = 37/60Now 1/5 part is filled in 1 hourHence 37/60 part will be filled = 5*37 /60
Arun
6 years ago
Thus 37/12 hours = 3 hours 5 min.Hence tank will be filled by 12:05 amHope it helpsConcept wiol be same.it might be that there is calculation mistake...but concept is important